OpenStudy (anonymous):

please explain how you would go about doing this.. find the sixth term of a geometric sequence with t(5)=24 and t(8)=3.

hartnn (hartnn):

know the formula for n'th term?

OpenStudy (anonymous):

that's the t(n)t(1)r(-1) right?

OpenStudy (anonymous):

but I don't know the first term, otherwise it's easy for me

hartnn (hartnn):

use that formula for t5 t5 = 24 n=5 \(24 =t_1r^4\) similarly use it for t(8) too

hartnn (hartnn):

\(\Large 8= t_1r^7\) do u get this?

OpenStudy (anonymous):

no I don't understand that at all. what is the r and how did you get the 7 and the 4 for it? is it just the term minus 1?

hartnn (hartnn):

\(t_n= t_1r^{n-1}\) for 8th term n=8 t8 is given as 3 \(\Large 3= t_1 r^{8-1}\) \(\Large 3= t_1 r^{7}\) now?

OpenStudy (anonymous):

ok yeah, how will you find the first term?

hartnn (hartnn):

\(21=t_1r^4 \\ 8=t_1r^7\) divide these 2 equations! and you will get 'r' (you need r also)

hartnn (hartnn):

24 not 21

OpenStudy (anonymous):

you mean divide 24 and 8?

hartnn (hartnn):

yep

OpenStudy (anonymous):

so 3. is that the first term or r?

hartnn (hartnn):

you divide both sides, isn't it ?